OpenStudy (anonymous):

A certain paint is to be mixed with 3 parts yellow to 2 parts blue. Twelve gallons of mixed paint are needed. How many gallons of blue must be used? Please Hurry!

OpenStudy (cathyangs):

The ratio of the paint recipe is 3 yellow : 2 blue. Adding 3 parts to 2 parts make 5 parts total. If you have 5 parts from this recipe, you can use proportions to find how much blue must be used: 2 parts blue / 5 total parts = X gallons blue / 12 total gallons. What do you get as your answer? (I cross-multipled to get 2* 12 = 5 * X, so 5X = 24, x = 24/5 gallons or 4 4/5.

OpenStudy (anonymous):

Oh my goodness, Thank you so much. your a life saver. @cathyangs